# Game-Theoretic Defenses for Robust Conformal Prediction Against Adversarial Attacks in Medical Imaging

### v1

@@ 5.2.1. Results for RQ1
Table 2. RQ1: Mean and Standard Deviation of Coverage, Size, SSCV, and Accuracy for OrganAMNIST
Attack Defensive Model Coverage (%) Size SSCV Accuracy (%)
FGSM FGSM 89.79 $\pm$ 0.48 1.47 $\pm$ 0.02 0.06 $\pm$ 0.01 80.93 $\pm$ 0.30
PGD 89.96 $\pm$ 0.49 1.46 $\pm$ 0.01 0.04 $\pm$ 0.01 83.34 $\pm$ 0.28
SPSA 90.01 $\pm$ 0.36 1.53 $\pm$ 0.01 0.06 $\pm$ 0.01 80.49 $\pm$ 0.15
Normal 89.96 $\pm$ 0.46 8.02 $\pm$ 0.05 0.13 $\pm$ 0.01 27.36 $\pm$ 0.24
PGD FGSM 89.86 $\pm$ 0.37 1.45 $\pm$ 0.01 0.05 $\pm$ 0.01 81.47 $\pm$ 0.20
PGD 89.93 $\pm$ 0.49 1.43 $\pm$ 0.01 0.04 $\pm$ 0.01 83.96 $\pm$ 0.25
SPSA 90.01 $\pm$ 0.34 1.51 $\pm$ 0.01 0.06 $\pm$ 0.01 80.77 $\pm$ 0.26
Normal 89.72 $\pm$ 0.36 8.91 $\pm$ 0.03 0.43 $\pm$ 0.03 14.86 $\pm$ 0.20
SPSA FGSM 89.88 $\pm$ 0.30 1.39 $\pm$ 0.01 0.06 $\pm$ 0.01 82.76 $\pm$ 0.18
PGD 89.92 $\pm$ 0.46 1.40 $\pm$ 0.02 0.04 $\pm$ 0.01 84.92 $\pm$ 0.22
SPSA 89.97 $\pm$ 0.45 1.41 $\pm$ 0.01 0.07 $\pm$ 0.01 82.61 $\pm$ 0.35
Normal 89.97 $\pm$ 0.29 7.72 $\pm$ 0.04 0.10 $\pm$ 0.01 31.70 $\pm$ 0.23
Table 3. RQ1: Mean and Standard Deviation of Coverage, Size, SSCV, and Accuracy for PathMNIST
Attack Defensive Model Coverage (%) Size SSCV Accuracy (%)
FGSM FGSM 90.19 $\pm$ 0.80 2.91 $\pm$ 0.06 0.05 $\pm$ 0.01 61.68 $\pm$ 0.56
PGD 90.22 $\pm$ 0.48 3.25 $\pm$ 0.03 0.04 $\pm$ 0.01 61.44 $\pm$ 0.29
SPSA 90.33 $\pm$ 0.56 2.92 $\pm$ 0.04 0.04 $\pm$ 0.01 58.04 $\pm$ 0.55
Normal 89.89 $\pm$ 0.79 5.86 $\pm$ 0.10 0.60 $\pm$ 0.02 22.28 $\pm$ 0.34
PGD FGSM 89.94 $\pm$ 0.39 2.95 $\pm$ 0.04 0.06 $\pm$ 0.01 61.60 $\pm$ 0.42
PGD 90.31 $\pm$ 0.43 3.20 $\pm$ 0.04 0.03 $\pm$ 0.01 62.20 $\pm$ 0.40
SPSA 90.25 $\pm$ 0.39 2.72 $\pm$ 0.03 0.04 $\pm$ 0.01 61.11 $\pm$ 0.58
Normal 91.00 $\pm$ 0.63 6.86 $\pm$ 0.07 0.72 $\pm$ 0.02 12.76 $\pm$ 0.10
SPSA FGSM 89.69 $\pm$ 0.61 2.97 $\pm$ 0.04 0.06 $\pm$ 0.01 59.45 $\pm$ 0.59
PGD 90.17 $\pm$ 0.46 3.05 $\pm$ 0.04 0.03 $\pm$ 0.01 64.42 $\pm$ 0.35
SPSA 90.03 $\pm$ 0.54 2.20 $\pm$ 0.04 0.03 $\pm$ 0.01 69.36 $\pm$ 0.51
Normal 89.82 $\pm$ 0.86 4.82 $\pm$ 0.14 0.39 $\pm$ 0.01 46.57 $\pm$ 0.62
Table 4. RQ1: Mean and Standard Deviation of Coverage, Size, SSCV, and Accuracy for TissueMNIST
Attack Defensive Model Coverage (%) Size SSCV Accuracy (%)
FGSM FGSM 90.03 $\pm$ 0.17 5.42 $\pm$ 0.02 0.19 $\pm$ 0.07 35.81 $\pm$ 0.16
PGD 90.03 $\pm$ 0.19 5.27 $\pm$ 0.02 0.21 $\pm$ 0.27 35.71 $\pm$ 0.11
SPSA 90.02 $\pm$ 0.20 6.50 $\pm$ 0.00 0.13 $\pm$ 0.10 35.82 $\pm$ 0.14
Normal 90.00 $\pm$ 0.21 5.91 $\pm$ 0.01 0.00 $\pm$ 0.00 14.91 $\pm$ 0.01
PGD FGSM 89.95 $\pm$ 0.30 5.40 $\pm$ 0.04 0.18 $\pm$ 0.06 35.98 $\pm$ 0.13
PGD 90.03 $\pm$ 0.19 5.24 $\pm$ 0.03 0.17 $\pm$ 0.26 35.83 $\pm$ 0.13
SPSA 90.07 $\pm$ 0.22 6.51 $\pm$ 0.01 0.13 $\pm$ 0.07 36.21 $\pm$ 0.19
Normal 90.06 $\pm$ 0.12 5.91 $\pm$ 0.01 0.00 $\pm$ 0.00 14.82 $\pm$ 0.02
SPSA FGSM 89.99 $\pm$ 0.28 5.40 $\pm$ 0.03 0.18 $\pm$ 0.05 36.04 $\pm$ 0.14
PGD 90.08 $\pm$ 0.20 5.25 $\pm$ 0.02 0.26 $\pm$ 0.34 35.83 $\pm$ 0.11
SPSA 89.98 $\pm$ 0.22 6.50 $\pm$ 0.01 0.17 $\pm$ 0.09 36.16 $\pm$ 0.14
Normal 89.99 $\pm$ 0.16 5.91 $\pm$ 0.00 0.00 $\pm$ 0.00 14.91 $\pm$ 0.02
5.2.2. Results for RQ2
In the second experiment, an unknown attack is applied to the test data, with the assumption that it belongs to one of the three predefined attack types: FGSM, PGD, or SPSA. We employ Algorithm 3 to determine the largest threshold across the calibration sets under different attacks, thereby constructing a conservative prediction set that ensures the coverage condition is met.
We evaluate the conformal prediction set’s performance in terms of coverage, size, and SSCV. The results are detailed in Tables 5, 6, and 7 for the respective datasets.
The findings from RQ2 indicate that the conservative approach effectively maintains the desired coverage level across all datasets, albeit with an increase in the average prediction set size due to the consideration of multiple potential attacks. In addition, the trade-off between coverage and prediction set size is evident, highlighting the balance between robustness and prediction precision.

### v2

@@ 5.2.1. Results for RQ1 and RQ2
Table 2. RQ1: Mean and Standard Deviation of Coverage, Size, and SSCV for OrganAMNIST
Attack Defensive Coverage (%) Size SSCV
Model APS RSCP VRCP-I VRCP-C APS RSCP VRCP-I VRCP-C APS RSCP VRCP-I VRCP-C
FGSM Normal 89.88 $\pm$ 0.51 89.40 $\pm$ 0.52 90.27 $\pm$ 0.50 91.70 $\pm$ 0.46 8.08 $\pm$ 0.05 8.13 $\pm$ 0.05 8.08 $\pm$ 0.05 8.39 $\pm$ 0.05 0.10 $\pm$ 0.03 0.10 $\pm$ 0.02 0.10 $\pm$ 0.03 0.10 $\pm$ 0.02
FGSM 89.82 $\pm$ 0.51 90.19 $\pm$ 0.50 89.45 $\pm$ 0.52 92.29 $\pm$ 0.45 1.48 $\pm$ 0.02 1.82 $\pm$ 0.02 1.36 $\pm$ 0.01 1.67 $\pm$ 0.02 0.06 $\pm$ 0.01 0.07 $\pm$ 0.02 0.08 $\pm$ 0.02 0.04 $\pm$ 0.00
PGD 90.07 $\pm$ 0.50 90.44 $\pm$ 0.49 90.19 $\pm$ 0.50 94.21 $\pm$ 0.39 1.43 $\pm$ 0.02 1.59 $\pm$ 0.02 1.25 $\pm$ 0.01 1.72 $\pm$ 0.02 0.05 $\pm$ 0.01 0.08 $\pm$ 0.02 0.09 $\pm$ 0.02 0.06 $\pm$ 0.01
SPSA 90.41 $\pm$ 0.49 90.83 $\pm$ 0.48 89.90 $\pm$ 0.51 92.21 $\pm$ 0.45 1.53 $\pm$ 0.02 2.16 $\pm$ 0.02 1.44 $\pm$ 0.01 1.71 $\pm$ 0.02 0.05 $\pm$ 0.01 0.06 $\pm$ 0.01 0.08 $\pm$ 0.02 0.04 $\pm$ 0.01
PGD Normal 90.47 $\pm$ 0.49 90.19 $\pm$ 0.50 89.99 $\pm$ 0.50 92.27 $\pm$ 0.45 8.90 $\pm$ 0.04 8.56 $\pm$ 0.05 8.89 $\pm$ 0.04 9.22 $\pm$ 0.04 0.53 $\pm$ 0.11 0.22 $\pm$ 0.04 0.54 $\pm$ 0.11 0.53 $\pm$ 0.12
FGSM 90.02 $\pm$ 0.50 90.30 $\pm$ 0.50 90.52 $\pm$ 0.49 93.14 $\pm$ 0.42 1.43 $\pm$ 0.02 1.80 $\pm$ 0.02 1.36 $\pm$ 0.01 1.67 $\pm$ 0.02 0.05 $\pm$ 0.01 0.08 $\pm$ 0.02 0.06 $\pm$ 0.01 0.04 $\pm$ 0.01
PGD 89.74 $\pm$ 0.51 90.04 $\pm$ 0.50 90.27 $\pm$ 0.50 93.79 $\pm$ 0.40 1.43 $\pm$ 0.02 1.54 $\pm$ 0.02 1.25 $\pm$ 0.01 1.70 $\pm$ 0.02 0.04 $\pm$ 0.01 0.09 $\pm$ 0.02 0.09 $\pm$ 0.03 0.05 $\pm$ 0.01
SPSA 90.35 $\pm$ 0.50 90.13 $\pm$ 0.50 90.47 $\pm$ 0.49 92.77 $\pm$ 0.43 1.50 $\pm$ 0.02 2.05 $\pm$ 0.02 1.41 $\pm$ 0.01 1.70 $\pm$ 0.02 0.06 $\pm$ 0.01 0.05 $\pm$ 0.01 0.07 $\pm$ 0.02 0.05 $\pm$ 0.01
SPSA Normal 89.15 $\pm$ 0.52 89.45 $\pm$ 0.52 89.31 $\pm$ 0.52 91.11 $\pm$ 0.48 7.69 $\pm$ 0.05 7.87 $\pm$ 0.05 7.70 $\pm$ 0.05 7.98 $\pm$ 0.05 0.10 $\pm$ 0.02 0.13 $\pm$ 0.03 0.10 $\pm$ 0.02 0.10 $\pm$ 0.02
FGSM 90.38 $\pm$ 0.49 90.21 $\pm$ 0.50 89.99 $\pm$ 0.50 93.73 $\pm$ 0.41 1.40 $\pm$ 0.02 1.66 $\pm$ 0.02 1.29 $\pm$ 0.01 1.61 $\pm$ 0.02 0.04 $\pm$ 0.01 0.07 $\pm$ 0.01 0.10 $\pm$ 0.02 0.05 $\pm$ 0.01
PGD 89.85 $\pm$ 0.51 90.27 $\pm$ 0.50 90.10 $\pm$ 0.50 93.90 $\pm$ 0.40 1.37 $\pm$ 0.01 1.47 $\pm$ 0.02 1.19 $\pm$ 0.01 1.64 $\pm$ 0.02 0.01 $\pm$ 0.00 0.09 $\pm$ 0.02 0.10 $\pm$ 0.03 0.05 $\pm$ 0.01
SPSA 90.35 $\pm$ 0.50 89.93 $\pm$ 0.50 91.20 $\pm$ 0.48 93.05 $\pm$ 0.43 1.41 $\pm$ 0.02 1.94 $\pm$ 0.02 1.34 $\pm$ 0.01 1.60 $\pm$ 0.02 0.07 $\pm$ 0.02 0.07 $\pm$ 0.01 0.06 $\pm$ 0.01 0.05 $\pm$ 0.01
CW Normal 89.74 $\pm$ 0.51 89.40 $\pm$ 0.52 89.65 $\pm$ 0.51 91.14 $\pm$ 0.48 7.74 $\pm$ 0.05 7.89 $\pm$ 0.05 7.74 $\pm$ 0.05 7.99 $\pm$ 0.05 0.19 $\pm$ 0.04 0.13 $\pm$ 0.02 0.19 $\pm$ 0.04 0.14 $\pm$ 0.03
FGSM 89.99 $\pm$ 0.50 90.52 $\pm$ 0.49 90.33 $\pm$ 0.50 93.00 $\pm$ 0.43 1.35 $\pm$ 0.01 1.69 $\pm$ 0.02 1.26 $\pm$ 0.01 1.57 $\pm$ 0.02 0.05 $\pm$ 0.01 0.06 $\pm$ 0.01 0.07 $\pm$ 0.02 0.04 $\pm$ 0.01
PGD 89.90 $\pm$ 0.51 90.80 $\pm$ 0.48 89.93 $\pm$ 0.50 94.52 $\pm$ 0.38 1.38 $\pm$ 0.01 1.48 $\pm$ 0.02 1.18 $\pm$ 0.01 1.66 $\pm$ 0.02 0.05 $\pm$ 0.01 0.10 $\pm$ 0.02 0.13 $\pm$ 0.03 0.06 $\pm$ 0.02
SPSA 90.55 $\pm$ 0.49 90.89 $\pm$ 0.48 90.80 $\pm$ 0.48 93.62 $\pm$ 0.41 1.40 $\pm$ 0.02 1.98 $\pm$ 0.02 1.32 $\pm$ 0.01 1.61 $\pm$ 0.02 0.06 $\pm$ 0.02 0.05 $\pm$ 0.01 0.08 $\pm$ 0.01 0.05 $\pm$ 0.01
Clean Normal 90.24 $\pm$ 0.50 91.84 $\pm$ 0.46 91.03 $\pm$ 0.48 93.17 $\pm$ 0.42 1.27 $\pm$ 0.01 1.26 $\pm$ 0.01 1.19 $\pm$ 0.01 1.44 $\pm$ 0.02 0.10 $\pm$ 0.02 0.07 $\pm$ 0.01 0.11 $\pm$ 0.02 0.05 $\pm$ 0.01
FGSM 90.72 $\pm$ 0.49 89.85 $\pm$ 0.51 89.82 $\pm$ 0.51 92.86 $\pm$ 0.43 1.29 $\pm$ 0.01 1.19 $\pm$ 0.01 1.17 $\pm$ 0.01 1.42 $\pm$ 0.02 0.09 $\pm$ 0.03 0.15 $\pm$ 0.04 0.15 $\pm$ 0.04 0.06 $\pm$ 0.01
PGD 90.97 $\pm$ 0.48 90.41 $\pm$ 0.49 90.86 $\pm$ 0.48 94.09 $\pm$ 0.40 1.24 $\pm$ 0.01 1.16 $\pm$ 0.01 1.17 $\pm$ 0.01 1.42 $\pm$ 0.02 0.04 $\pm$ 0.01 0.09 $\pm$ 0.03 0.08 $\pm$ 0.02 0.05 $\pm$ 0.01
SPSA 90.49 $\pm$ 0.49 89.85 $\pm$ 0.51 90.49 $\pm$ 0.49 93.81 $\pm$ 0.40 1.26 $\pm$ 0.01 1.19 $\pm$ 0.01 1.18 $\pm$ 0.01 1.42 $\pm$ 0.02 0.07 $\pm$ 0.02 0.11 $\pm$ 0.03 0.13 $\pm$ 0.03 0.05 $\pm$ 0.01